1.For a table manufacturing company, selling price for a table is $200.00 per Unit, Variable cost is $125.00 per Unit, rent is $2,797.00 per month and insurance is $1,396.00 per month. How much is the company contributing to covering their fixed costs or generating profits?

2.

You run a school in Florida. Fixed monthly cost is $5,378.00 for rent and utilities, $5,514.00 is spent in salaries and $1,265.00 in insurance. Also every student adds up to $100.00 per month in stationary, food etc. You charge $674.00 per month from every student now.

You are considering moving the school to another neighborhood where the rent and utilities will increase to $10,911.00, salaries to $6,326.00 and insurance to $2,465.00 per month. Variable cost per student will increase up to $185.00 per month. However you can charge $1,029.00 per student. At what point will you be indifferent between your current mode of operation and the new option?

3.A restaurant sells pizza at a rate of $10.58/slice. Expenses for the restaurant include raw material for pizza at $5.58 per slice, $167.00 as monthly rental and $65.00 monthly as insurance. How many slices should the restaurant sell in a month to break even?
Answer #1

Ans 1.  As per the Question , there are two Fixed cost i.e Rent $ 2797.00 per month and Insurance $1396.00 per month.

Total Fixed Cost is $4193.00

Selling Price per unit is $200

Variable Cost per unit is $125

Contribution = Selling price - Variable Cost

Contribution = $200 - $125

Contribution = $75 per unit

Company contributing $75 per unit to cover their Fixed Cost and to generate profits.

Ans 2.

Current Position :

Particulars Cost ($)
Rent and Utilities 5378
Salaries 5514
Insurance 1265
Total Fixed Cost 12157

Variable Cost is $100 per student per month

Fee per student per month is $ 674

Let say Number of students is "n".

Profit = Selling Price X Units - Variable Cost X Units - Fixed Cost

Profit = 674n - 100n - 12157

Alternatively if School is moved to next building

Particulars Cost ($)
Rent and Utilities 10911
Salaries 6326
Insurance 2465
Total Fixed Cost 19702

Total Variable Cost is $185 per student per month

Fee per student per month $ 1029.

Profit = 1029n - 185n - 19702

The point at which there is will be no indifference is when we have profits same in both situations.

674n - 100n - 12157 = 1029n - 185n - 19702

574n -12157 = 844n - 19702

19702-12157 = 844n - 574n

7545 = 270n

270n = 7545

n = 27.944

i.e When 28 students were there will no difference in current or in proposed neighborhood.

If Number of students is less than 28 then Current position is better

If Number of Students is more than 28 then Proposed neighborhood is better.

Ans.3.

Break-Even Point (Units) = Fixed Costs ÷ (Revenue per Unit – Variable Cost per Unit)

Fixed Cost are Rental and Insurance i.e $167 and $65 respectively. Total Fixed Cost is $232.

Total Variable Cost per slice is $ 5.58.

Selling Price $10.58 per slice.

Break Even Point (Units) = 232 ÷ (10.58-5.58)

Break Even Point (Units) = 232 ÷ 5

Break Even Point (Units) = 46.4 Slices

To Breakeven restaurant has to sell 47 Slices .